OpenStudy (anonymous):

One number is 5 more than another number. Find these numbers if their product is -3.

OpenStudy (anonymous):

I just need to know how to set up the problem :)

myininaya (myininaya):

lets call one number x and the other y one number is 5 more than another number x = 5 + y so we have one equation x=5+y and we know x and y has a product of -3=>xy=-3 so the other equation is xy=-3 we know x=5+y so plug this x into the one that is in xy=-3 so we have (5+y)y=-3

OpenStudy (anonymous):

to setup: x= 5+y (one number is 5 more than the othe rnumber) xy = -3 (their product is -3)

OpenStudy (anonymous):

Did you get -5+_ sqrt13/2??

myininaya (myininaya):

(5+y)y=-3 5y+y^2=-3 y^2+5y=-3 y^2+5y+(5/2)^2=-3+(5/2)^2 (y+5/2)^2=-3+25/4 (y+5/2)^2=-3*4/4+25/4 (y+5/2)^2=-12/4+25/4 (y+5/2)^2=13/4 take square root of both sides y+5/2=plus or minus sqrt{13/4} y+5/2=plus or minus sqrt{13}/2 y=-5/2 plus or minus sqrt{13}/2 so we have y=-5/2 + sqrt{13}/2 or y=-5/2 - sqrt{13}/2

myininaya (myininaya):

and so x=5+y so just plug your two y's into there to find your x's

OpenStudy (anonymous):

Thank you myininaya and laksmanv :)

myininaya (myininaya):

np crazy how they didn't choose a word problem that worked out more nicely